# FADE: A Dataset for Detecting Falling Objects around Buildings in Video

### v1

@@ Abstract
Falling objects from buildings can cause severe injuries to pedestrians due to the great impact force they exert. Although surveillance cameras are installed around some buildings, it is challenging for humans to capture such events in surveillance videos due to the small size and fast motion of falling objects, as well as the complex background. Therefore, it is necessary to develop methods to automatically detect falling objects around buildings in surveillance videos. To facilitate the investigation of falling object detection, we propose a large, diverse video dataset called FADE (FAlling Object DEtection around Buildings) for the first time. FADE contains 1,881 videos from 18 scenes, featuring 8 falling object categories, 4 weather conditions, and 4 video resolutions. Additionally, we develop a new object detection method called FADE-Net, which effectively leverages motion information and produces small-sized but high-quality proposals for detecting falling objects around buildings. Importantly, our method is extensively evaluated and analyzed by comparing it with the previous approaches used for generic object detection, video object detection, and moving object detection on the FADE dataset. Experimental results show that the proposed FADE-Net significantly outperforms other methods, providing an effective baseline for future research. The dataset and code are publicly available at https://fadedataset.github.io/FADE.github.io/.
@@ I Introduction
Over the past few decades, with the continuous expansion of urbanization, high-rise buildings have sprung up. Some people, who throw objects from tall buildings arbitrarily, causing injuries and incidents one after another. According to a report [1] of the U.S. Bureau of Labor Statistics, there are more than 50,000 “struck by falling objects” injuries every year in USA. We use a specific example to illustrate the possible hazards of objects falling from tall buildings. If a 200-gram apple drops from a 30-meter building, the impact duration of the apple is approximate 0.01 seconds. Based on the momentum theorem [2], it means that the ground will bear an impact of about 49.5-kilogram weight. As depicted in Figure 1, this type of incident profoundly endangers public safety. To reduce such incidents, some countries have legislated to rule that throwing objects from buildings is illegal, e.g., USA [3], Singapore [4], China [5], etc. With more and more cameras installed around buildings, it is necessary to explore how to effectively detect these falling object events by utilizing the surveillance videos.
Figure 1: The incident of falling object around the building produces great danger and harm to human life and social security.
@@
Intelligent video surveillance (IVS) has become a crucial technology for public safety[6]. Recent advances in computer vision promoted the development of IVS systems for various security applications, including object detection [7, 8, 9], anomaly detection[10, 11, 12, 13], human identification [14, 15, 16], tracking[17, 18, 19], and behavior understanding [20, 21, 22, 23]. These computer vision based methods have advantages of low cost, high precision, and manpower saving. Inspired by the successful applications of these computer vision techniques, some research institutions and government departments have begun to explore computer vision algorithms for detecting the incidents of falling objects around buildings. The task of falling object detection around buildings (FODB) can be considered as a special case of moving object detection (MOD). Although there have been several MOD datasets, such as SABS [24], CDnet 2014 [25], GTFD [26], and LASIESTA [27], the objects of interest in these datasets differ from those in FODB largely. The falling objects in FODB are usually much smaller and move faster, making them more difficult to detect.
A comparison between the FODB task and the MOD task is shown in Figure 2. Where in an existing MOD dataset, the moving object occupies a large area of the image. In contrast, the falling object from the building takes up a small area of the image, causing it is difficult to distinguish the falling objects from the background. Generally, the motion of the falling object in the FODB task is fast, resulting in 1) motion blur and 2) a large displacement of the object between adjacent video frames, which are uncommon in the MOD dataset. Furthermore, none of the existing MOD datasets cover the categories of falling objects around buildings. Consequently, it is necessary to collect a large diverse dataset as well as propose an effective baseline method for FODB. Such efforts are crucial for advancing the progress and practical application of FODB.
@@
In this work, for the first time, we present a completely new video dataset termed FADE for FODB, which consists of 1,881 videos under various scenes. Specifically, the dataset contains 164,314 video frames in total, all of which are annotated. Our dataset is diverse, covering 18 scenes, 8 object categories, 4 video resolutions, 4 weather conditions, 3 camera angles, and 2 lighting conditions. As can be seen in Table I, our dataset FADE has richer data compared to the existing MOD datasets. The size of the falling object in our FADE is small (shown in Table II) with a median area of about 20 pixels, which is much smaller than the small object ($area<32^{2}$ pixels) defined in the COCO dataset [28]. We also propose a new baseline method, FADE-Net, which utilizes motion cues to complement appearance features and modifies the region proposal network to generate high-quality proposals for small falling objects.
We use the evaluation metrics in the MOD task to benchmark several methods of 3 different tasks (MOD, generic object detection, and video object detection) on our dataset. We also explore a new metric, time range overlap (TRO), to evaluate the performance of the detection methods on localizing the object falling incidents. The experimental results indicate that FODB is a challenging task and validate the effectiveness of our presented method.
@@
  • We construct a new video dataset called FADE which is the first dataset used for falling object detection around buildings (FODB). This dataset is large and diverse, which covers various scenes and complex conditions.
  • We explore a new baseline method FADE-Net for the FODB task, which effectively utilizes motion cues and generates high-quality proposals for small-sized falling objects detection.
@@ II Related Work
Moving Object Detection Dataset. MOD is a basic task in computer vision, and some MOD datasets have been released for training and testing the MOD algorithms. Early MOD datasets, e.g. Wallflower [29], SegTrack [30], and I2R [31], are small in scale. Wallflower [29] is one of the earliest MOD datasets, which has solely 7 video sequences, and each sequence has only 1 ground truth (GT) frame. SegTrack [30] consists of only 6 video sequences and 215 annotated frames. 10 video sequences are provided in I2R [31], including videos with dynamic backgrounds, challenging weather conditions, and gradual illumination variations. Later, some large scale and complicated MOD datasets are constructed. For instance, SABS [24] contains videos with 10 categories, and each video has 800 training frames. Some testing video frames in FBMS 59 [32] and SegTrack V2 [33] contain multiple moving objects. GTFD [26] provides a collection of 25 videos with both rigid and non-rigid moving objects. LASIESTA [27] consists of 45 videos and 18,425 video frames, recorded by the moving and static cameras. CDnet 2012 [34] and CDnet 2014 [25] are presented as benchmarks in IEEE Change Detection Workshop. CDnet 2014 [25] increases 22 videos and 5 categories on the basis CDnet 2012 [25]. Captured in the outdoor environment, BMC 2012 [35] includes a total of 29 real and synthetic videos in different weather conditions. DAVIS 2016 [36] supplies 50 high-quality and densely annotated videos. SBI [37] composes of 14 image sequences and the corresponding GT backgrounds, which is the first dataset used to evaluate the background initialization approaches.
@@
Unlike these MOD datasets, FADE is the first dataset for FODB which contains numerous videos and diverse scenes. The comparison between our FADE and the prior MOD datasets is shown in Table I.
TABLE I: Comparison of our FADE dataset with the prior MOD datasets. “GT frames” means the Ground Truth frames. The datasets are sorted by the publishing time in ascending order. The maximum quantity and the second maximum quantity are highlighted in bold and underline, respectively. Our FADE dataset has the largest number of videos, total frames, GT frames, and scenes.
Dataset Total videos $\uparrow$ Total frames GT frames Scenes
SABS [24] 9 6,408 2,000 9
SegTrack V2 [33] 14 976 976 14
SBI [37] 14 5,024 14 14
GTFD [26] 25 1,067 1,067 7
CDnet 2014 [25] 31 90,000 90,000 11
LASIESTA [27] 48 18,425 18,425 4
DAVIS 2016 [36] 50 3,455 3,455 15
FADE (Ours) 1,881 164,314 164,314 18
@@
Moving Object Detection. MOD has been extensively investigated due to its wide range of applications [38, 39, 40]. Many unsupervised MOD methods have been explored, which can be classified into two categories, i.e. background modeling and feature extraction. Specifically, for background modeling, parametric based Gaussian mixture methods [9, 41, 42] are proposed to model the background in MOD. Barnich et al. [8] update the background by applying a novel random selection strategy. Baf et al. [43] apply Choquet integral [44] as an aggregation operator to aggregate color and texture features for MOD. Lin et al. [45] present a dual-rate background modeling framework to detect foreground objects, utilizing both short-term and long-term background models to improve the accuracy of foreground inference. For feature extraction, Yang et al. [46] adopt an optical flow based method to detect moving objects. Zhou et al. [11] use motion information to enhance detection with motion fusion blocks that compressing video clips into a single image. Thanikasalam et al. [47] exploit a target-specific Siamese attention network that employs residual and channel attention modules to capture the global and channel-wise information of moving objects. Shang et al. [48, 49] consider MOD as a robust principal component analysis problem involving robust subspace learning and tracking [50, 51, 52, 53]. To address sudden and gradual background changes, Dong et al. [54] explore a clustering feature space method to represent different background appearances. The developments of deep learning [55, 56, 57, 58, 59] have greatly promoted the progress of supervised MOD methods. Deep learning-based MOD approaches can be divided into six main categories: basic CNN [60, 61, 62, 63, 64], deep CNN [65, 66], 3D CNN [67, 68], ConvLSTM [69], FCN [70], and GAN [71, 72]. Although numerous MOD methods have been proposed, few can be applied to FODB. We select representative state-of-the-art MOD methods to benchmark our dataset.
Falling Object Detection. There are a few works concentrated on falling object detection. Some use different sensors to handle this task. 3D point cloud of the object generated by the distance sensor, is applied to predict the falling probability of the object in the indoor or at the building site [73]. Alvare et al. [74] utilize the ultrasonic sensor and the signal coding sequence technology, which is based on complementary coding, to detect falling object on railway tracks. By employing the object detection method, Yang et al. [75] obtain the positional relationship between the objects with falling risk and the workers at the building site, and establish an automatic management system for evaluating the falling risk. Different from these methods, our approach utilizes motion cues to complement appearance features for better falling object detection around buildings.
@@ III The Proposed FADE Dataset
We collect and release the first video dataset for falling object detection around buildings publicly to promote its development. In this section, we introduce our FADE dataset in detail from the aspects of metadata, dataset construction, dataset division and statistics, dataset ethics, license, maintenance plan, and evaluation metrics.
@@ III-A Metadata
To provide a comprehensive FODB dataset, we collect videos in diverse weather conditions, lighting cases, scenes, camera angles, and video resolutions. Example video frames are shown in our dataset (https://fadedataset.github.io/FADE.github.io/). Our metadata is defined as follows.
@@
Weather Condition. Different weather conditions lead to different light intensities, which affect the contrast between falling object and the background. Weather condition is one of the most important influences. We design our dataset with 4 weather conditions: fair, cloudy, overcast, and rainy, which can be seen in Figure 3.
Figure 3: Four weather conditions in our FODB dataset including fair, cloudy, overcast and rainy.
Figure 4: We provide videos of two modes (RGB mode and grayscale mode) in various scenes, such as dormitories, classroom buildings, office buildings, etc.
Lighting Condition. Generally, when the light intensity is lower than 0.04 Lux, the image of the surveillance video will change from the RGB mode to the grayscale mode. Both the RGB mode and the grayscale mode videos are collected in various scenes to enrich our dataset. As shown in Figure 4, we provide videos in both the RGB mode and the grayscale mode with light intensity more and less than 0.04 Lux, respectively.
Scene. The scene of the falling incident is diverse. Our FADE dataset includes 18 different scenes, covering almost all the scenarios where the falling incidents may occur. E.g., classroom buildings, office buildings, dormitories, apartments, and buildings under construction (shown in Figure 4). In our dataset, the scene indicates one viewpoint of a building. In different scenes, the buildings appear different as they are captured in different viewpoints.
Camera Angle. The surveillance cameras for FODB are normally installed 30 meters away from the high-rise buildings, and different floors of the buildings are monitored by cameras from different angles. In our dataset, we consider three different camera angles: 30°, 45°, and 60°.
@@ III-B Dataset construction
This section introduces how we build our dataset in detail. Our dataset documentation, annotation instruction, intended utilization, and structured metadata are published at the website https://fadedataset.github.io/FADE.github.io/index.html. The example videos and evaluation code are provided in our website. Our data does not contain any personally identifiable information or offensive content. We bear all responsibility in case of violation of rights and confirmation of the data license.
Data preparation. The diversity of object categories is important for training models for falling object detection around building. As we have stated above, the falling objects in our dataset cover a number of common classes, including clothes, shoes, kitchen waste, books, spitballs, bottles, packaging bags, and packaging boxes. Each object category also includes a rich variety of object instances. For example, the category of kitchen waste includes banana peels, uneaten apples, etc. Besides, light intensity varies with weather condition changes, which would affect the detection performance of falling objects. In addition, the appearance difference between the falling object and the background in a cloudy scene is smaller than that in a fair scene, and raindrops in rainy scenes will also affect the detection accuracy.
Data Collection. We recruit seven volunteers and spend a year and a half to collect the video data. We throw objects from the high-rise buildings of the corresponding university, and use a wide dynamic range camera, equipped with a $1/3^{{}^{\prime\prime}}$ progressive scanning CMOS sensor and dot matrix LED infrared lamp to record the whole process of these events. The highest output video quality of the camera is $2592\times 1520$ $30$ FPS. The motion blur caused by the high-speed moving objects can be reduced if using a global-shutter and high-FPS CMOS. However, this expensive CMOS produces the KTC noise [76]. Besides, it generates more video frames per second, leading to a higher storage cost. More importantly, the already installed cameras around the world for monitoring the falling object are mostly equipped with the general economic sensors. We therefore do not use the camera with global-shutter and high-FPS CMOS to capture video data.
@@
Data Annotation. The annotation process lasts for half a year, including two rounds. In the first round, the novice annotator labels the video. In the second round, the expert annotator checks the annotation to improve the quality. Different from the pixel-by-pixel annotation method of the MOD datasets [24, 26, 37], we adopt the annotation manner used in the object detection task to generate the bounding box around each falling object.

### v3

@@ Abstract
Objects falling from buildings, a frequently occurring event in daily life, can cause severe injuries to pedestrians due to the high impact force they exert. Surveillance cameras are often installed around buildings to detect falling objects, but such detection remains challenging due to the small size and fast motion of the objects. Moreover, the field of falling object detection around buildings (FODB) lacks a large-scale dataset for training learning-based detection methods and for standardized evaluation. To address these challenges, we propose a large and diverse video benchmark dataset named FADE. Specifically, FADE contains 2,611 videos from 25 scenes, featuring 8 falling object categories, 4 weather conditions, and 4 video resolutions. Additionally, we develop a novel detection method for FODB that effectively leverages motion information and generates small-sized yet high-quality detection proposals. The efficacy of our method is evaluated on the proposed FADE dataset by comparing it with state-of-the-art approaches in generic object detection, video object detection, and moving object detection. The dataset and code are publicly available at https://fadedataset.github.io/FADE.github.io/.
@@ I Introduction
Over the past few decades, with the continuous expansion of urbanization, high-rise buildings have sprung up. Some residents of these buildings throw objects from above without caution, potentially leading to repeated injuries and incidents. According to a report [1] of the U.S. Bureau of Labor Statistics, there are more than 50,000 “struck by falling objects” injuries every year in USA. To highlight the potential danger posed by objects falling from tall buildings, we present a specific example: If a 200-gram apple falls from a 30-meter-high building, the impact duration is approximately 0.01 seconds. Based on the momentum theorem [2], this results in an equivalent impact force of roughly 49.5 kilograms. Such incidents can pose serious threats to public safety (see Figure 1 for visualization). To mitigate such incidents, several countries have enacted laws prohibiting the act of throwing objects from buildings, including the USA [3], Singapore [4], and China [5].
Figure 1: Falling object incidents around buildings may pose serious risks to human life and public safety.
@@
At the same time, intelligent video surveillance (IVS) has become a critical technology for ensuring public safety [6], fueled by recent advances in computer vision, including object detection [7, 8, 9], anomaly detection [10, 11, 12, 13], human identification [14, 15, 16, 17], tracking [18, 19, 20, 21], and video understanding [22, 23, 24, 25, 26]. These computer vision-based IVS methods provide benefits such as low cost, high accuracy, and reduced dependence on manual labor. Inspired by the successful applications of these techniques and supported by the availability of surveillance cameras around buildings, several research institutions and government agencies have begun exploring IVS algorithms for detecting falling object incidents around buildings (FODB). However, these IVS technologies often rely on large-scale datasets to train learning-based models, yet such a dataset is currently unavailable in the FODB domain. In fact, the FODB task can be regarded as a special case of the moving object detection (MOD) task, and it may appear feasible to utilize existing MOD datasets, such as SABS [27], CDnet 2014 [28], GTFD [29], and LASIESTA [30], to train FODB algorithms. However, this approach is limited by the significant disparity between the two tasks. Specifically, moving (falling) objects in FODB are typically much smaller and move at higher speeds than the objects in standard MOD scenarios.
To better illustrate the differences between the FODB task and the MOD task, we present a visual comparison in Figure 2. As shown in Figure 2, moving objects in existing MOD datasets typically occupy large areas of the image, whereas falling objects in the FODB task are much smaller and cover only a small portion of the scene. Besides, falling objects exhibit rapid motion, resulting in (1) motion blur and (2) large displacements between adjacent video frames. These factors make falling objects difficult to detect and are rarely encountered in MOD datasets. Furthermore, none of the existing MOD datasets include categories corresponding to falling objects around buildings. Consequently, it is necessary to construct a large and diverse benchmark dataset to evaluate the performance of FODB algorithms. Such efforts are critical for advancing research and facilitating real-world deployment in FODB domain.
@@
TABLE I: Comparison of our FADE dataset with the prior MOD datasets. “GT frames” means the Ground Truth frames. The datasets are sorted by the publishing time in ascending order. As can be seen, our FADE dataset has the largest number of videos, total frames, GT frames, and scenes.
Dataset Total videos $\uparrow$ Total frames GT frames Scenes Frame-wise IOU (%)
SABS [27] 9 6,408 2,000 9 43
SegTrack V2 [31] 14 976 976 14 27
SBI [32] 14 5,024 14 14 31
GTFD [29] 25 1,067 1,067 7 37
CDnet 2014 [28] 31 90,000 90,000 11 33
LASIESTA [30] 48 18,425 18,425 4 35
DAVIS 2016 [33] 50 3,455 3,455 15 32
FADE (Ours) 2,611 245,177 245,177 25 13
In this work, we present a video benchmark dataset for FODB, termed FADE, which comprises 2,611 videos captured across diverse scenes. Specifically, the dataset includes 245,177 annotated video frames spanning 25 scenes, 8 object categories, 4 video resolutions, 4 weather conditions, 3 camera angles, and 2 lighting conditions. As can be seen in Table I, our dataset FADE has richer data compared to the existing MOD datasets. The size of the falling object in our FADE is small with a median area of about 20 pixels, which is also much smaller than the small object ($area<32^{2}$ pixels) defined in the COCO dataset [34]. We also introduce a new baseline method, FADE-Net, which leverages motion cues to complement appearance features and incorporates a small-object mining region proposal network to generate high-quality proposals for small falling objects. We use the evaluation metrics in the MOD task to benchmark several methods of 3 different tasks (MOD, generic object detection, and video object detection) on our dataset. We also explore a new metric, time range overlap (TRO), to evaluate the performance of the detection methods on localizing the object falling incidents. The experimental results indicate that FODB is a challenging task and validate the effectiveness of our presented method.
@@
  • We construct a new video dataset called FADE, which, in terms of application scenarios, is the first dataset for falling object detection around buildings (FODB). This dataset is large and diverse, which covers various scenes and complex conditions.
  • We explore a new baseline method FADE-Net for the FODB task, which effectively utilizes motion cues and can generate high-quality proposals for small-sized falling objects detection.
@@ II-A Moving Object Detection Dataset
Moving Object Detection (MOD) is a basic task in computer vision, and some MOD datasets have been released for training and testing the MOD algorithms. Early MOD datasets, such as Wallflower [35], SegTrack [36], and I2R [37], are limited in scale. For instance, Wallflower [35] is one of the earliest MOD datasets, containing only 7 video sequences, each with a single ground truth frame. SegTrack [36] consists of only 6 video sequences and 215 annotated frames. I2R [37] provides 10 video sequences, including scenes with dynamic backgrounds, challenging weather conditions, and gradual illumination variations. Later, some large scale and complicated MOD datasets are constructed. For instance, SABS [27] contains videos with 10 categories, and each video has 800 training frames. Some testing video frames in FBMS 59 [38] and SegTrack V2 [31] contain multiple moving objects. GTFD [29] provides a collection of 25 videos with both rigid and non-rigid moving objects. LASIESTA [30] consists of 45 videos and 18,425 video frames, recorded by the moving and static cameras. CDnet 2012 [39] and CDnet 2014 [28] serve as benchmarks for the IEEE Change Detection Workshop, with CDnet 2014 adding 22 videos and 5 new categories to CDnet 2012. Captured in the outdoor environment, BMC 2012 [40] includes a total of 29 real and synthetic videos in different weather conditions. DAVIS 2016 [33] supplies 50 high-quality and densely annotated videos. SBI [32] comprises 14 image sequences along with corresponding ground truth backgrounds and is the first dataset designed for evaluating background initialization MOD methods.
@@
Unlike these MOD datasets, FADE is the first dataset for the falling object detection around buildings task which contains numerous videos and diverse scenes. The quantitative comparison between our FADE and the prior MOD datasets is shown in Table I.
II-B Moving Object Detection
@@
MOD has been extensively investigated due to its wide range of applications [41, 42, 43]. Many unsupervised MOD methods have been explored and can be broadly classified into two categories: background modeling and feature extraction. Specifically, for background modeling, parametric Gaussian mixture methods [9, 44, 45] have been proposed to represent the background in MOD. Barnich et al. [8] update the background by applying a novel random selection strategy. Baf et al. [46] apply Choquet integral [47] as an aggregation operator to aggregate color and texture features for MOD. Lin et al. [48] propose a dual-rate background modeling framework for foreground object detection, which leverages both short-term and long-term background models to enhance the accuracy of foreground inference. For feature extraction, Yang et al. [49] adopt an optical flow based method to detect moving objects. Zhou et al. [11] use motion information to enhance detection with motion fusion blocks that compressing video clips into a single image. Thanikasalam et al. [50] exploit a target-specific Siamese attention network that employs residual and channel attention modules to capture the global and channel-wise information of moving objects. Shang et al. [51, 52] consider MOD as a robust principal component analysis problem involving robust subspace learning and tracking. To address sudden and gradual background changes, Dong et al. [53] explore a clustering feature space method to represent different background appearances.
The developments of deep learning [54, 55, 56] have greatly promoted the progress of supervised MOD methods. Early approaches can be broadly classified into six main categories: basic CNN [57, 58, 59, 60, 61], deep CNN [62, 63], 3D CNN [64, 65], ConvLSTM [66], FCN [67], and GAN [68, 69]. Specifically, CTFU-Net [70] hierarchically integrates the local feature extraction capabilities of CNNs with the global context modeling of Transformers to address dynamic scenes. TransBlast [71] employs an SVD-based subspace loss and Barlow Twins self-supervision to preserve foreground details while reducing the need for extensive annotations. GraphMOS-U [72] and GraphIMOS [73] are representative GNN-based approaches. GraphMOS-U [72] enables minimal-annotation underwater MOD by initializing Mask R-CNN with domain-specific feature fusion and Sobolev optimization. GraphIMOS [73] replaces transductive graphs with inductive block-diagonal GNNs to support real-time deployment on unseen videos.
Although numerous MOD methods have been proposed, few can be applied to the FODB task. In this work, we provide benchmark results of state-of-the-art MOD methods on our proposed FODB benchmark dataset.
II-C Falling Object Detection
There are a few works [74, 75, 76] concentrated on falling object detection. Specifically, [74] targets the prediction of potential falling objects (much like moving objects) in indoor environments by leveraging 3D point cloud data captured by distance sensors. [75] focuses on detecting objects which have already fallen (these are stationary) on railway tracks using ultrasonic sensors and signal coding sequence technology. [76] addresses falling hazards, that is, identifying stationary places a person might fall from or into, by analyzing the positional relationship between hazardous objects and workers at construction sites.
FODB is a critical task, as such incidents occur frequently, with approximately 50,000 cases reported annually in the United States [1]. These incidents pose fatal risks to pedestrians due to the high impact force of falling objects (see Section I). However, existing falling object detection methods [74, 75, 76] are not designed with small object detection and motion information utilization in mind, making them unsuitable for direct application to FODB scenarios, which involve detecting fast-moving small objects. Besides, the FODB field also lacks a public dataset, limiting both the training of learning-based methods and the evaluation of falling object detectors. To solve these issues, we construct the first large-scale benchmark dataset for FODB, and introduce a dedicated FODB method, FADE-Net.
@@ III The Proposed FADE Dataset
To advance research in the area of falling object detection around buildings (FODB), we construct and release the first benchmark dataset. In this section, we provide a detailed introduction to our FADE dataset, covering metadata, dataset construction, dataset splits and statistics, ethical considerations, licensing, maintenance plan, and evaluation metrics.
Figure 3: To make our dataset design more accessible, we follow the commonly used Kaggle weather dataset [77] and categorize weather conditions into four broad types: fair, cloudy, overcast, and rainy.
@@ III-A Metadata
To provide a comprehensive FODB dataset, we collect videos in diverse weather conditions, lighting cases, scenes, camera angles, and video resolutions. Example video frames are shown in our dataset (https://fadedataset.github.io/FADE.github.io/). Our metadata is defined as follows:
@@
Weather Condition. Different weather conditions lead to different light intensities, which affect the contrast between falling object and the background. To make our dataset design more accessible, we follow the commonly used Kaggle weather dataset [77] and classify weather conditions into four broad categories: fair, cloudy, overcast, and rainy, as illustrated in Figure 3.
Figure 4: Our dataset includes two modes (RGB and grayscale), captured across various scenes such as dormitories, classroom buildings, and office buildings.
Lighting Condition. Generally, when the light intensity is lower than 0.04 Lux, the image of the surveillance video will change from the RGB mode to the grayscale mode. Therefore, to ensure the generality of our dataset, we provide videos in both RGB mode and grayscale mode, corresponding to light intensities greater than and less than 0.04 Lux, respectively, as shown in Figure 4.
Scene. The scenes in our dataset are diverse. Specifically, the FADE dataset includes 18 distinct scenes that cover a wide range of environments where falling incidents may occur, such as classroom buildings, office buildings, dormitories, apartments, and buildings under construction, as illustrated in Figure 4. Notably, in our dataset, a scene refers to a specific viewpoint of a building. Across different scenes, the buildings appear distinct due to variations in the camera viewpoints.
Camera Angle. The videos in our dataset cover 3 different camera angles: 30°, 45°, and 60°. This design reflects real-world FODB surveillance setups, where cameras are typically installed 30 meters away from buildings, with different floors monitored using cameras positioned at varying angles.
@@ III-B Dataset construction
This section provides a detailed description of our dataset construction process, including data preparation, data collection, data format, and data annotation. Moreover, we provide dataset documentation, annotation guidelines, intended use cases, structured metadata, example videos, and evaluation code on our website https://fadedataset.github.io/FADE.github.io/index.html.
Data Preparation. The diversity of object categories is important for training models for falling object detection around building. As stated above, the falling objects in our dataset span several common categories, including clothes, shoes, kitchen waste, books, spitballs, bottles, packaging bags, and packaging boxes. In addition, each category contains a diverse set of object instances. For example, the kitchen waste category includes items such as banana peels and uneaten apples.
Data Collection. We recruit seven volunteers and spend a year and a half to collect the video data. Specifically, we throw objects from the high-rise buildings of the corresponding university, and use a wide dynamic range camera, equipped with a $1/3^{{}^{\prime\prime}}$ progressive scanning CMOS sensor and dot matrix LED infrared lamp to record the whole process of these events. The highest output video quality of the camera is $2592\times 1520$ @ $30$ FPS. We observe that some falling objects in our recorded video exhibit motion blur due to their high falling speed. Although recording videos with CMOS cameras equipped with global shutters and high FPS can mitigate this issue, the use of such specialized cameras presents two main drawbacks: (1) videos captured with these cameras often suffer from KTC noise [78], introducing visual artifacts; and (2) most existing surveillance systems deployed worldwide for monitoring falling objects use general-purpose, low-cost sensors. As a result, models trained on data from these specialized CMOS cameras may be less applicable to real-world FODB scenarios.
@@
Data Annotation. The annotation process of our dataset lasts for half a year, including two rounds. In the first round, the novice annotator labels the video. In the second round, the expert annotator checks the annotation to improve the quality. Different from the pixel-by-pixel annotation method of the MOD datasets [27, 29, 32], we adopt the annotation manner used in the object detection task to generate the bounding box around each falling object.